# Quake: Adaptive Indexing for Vector Search

### v1

@@ 7.6 Comparison with Early Termination Methods
Table 5 compares early-termination methods on Sift1M, highlighting the tradeoff between query latency, tuning time, and recall, we do not include ground truth generation time in the tuning time for the baselines. APS analytically estimates recall at query time, eliminating offline tuning entirely, while achieving latency within 20% of the oracle across all recall targets. Fixed selects a static $nprobe$ per target via an expensive offline binary search (up to 424s), and SPANN similarly performs a binary search and tunes a centroid-distance threshold; both closely match recall targets but incur higher latency at 90% and 99% recall. LAET trains a per-query prediction model, incurring moderate tuning overhead (81–232s), and matches recall targets with slightly higher latency compared to APS. Auncel is the most similar method to APS, as it aims to analytically estimate recall using partition intersection volumes, however its volume estimation requires calibration, and it is a conservative method, overshooting recall. We tune Auncel by binary searching a geometric parameter ($a$), overshooting recall significantly (up to 8.1 pp) and increasing latency by up to 69% compared to APS. Finally, the Oracle, which scans the minimal amount of partitions per-query, serves as a practical lower bound on achievable latency, though with prohibitively high tuning cost. The tuning overhead of the baselines demonstrates a significant burden for online scenarios where the queries, data, and index change. The tuning burden worsens at scale, where ground truth generation cost grows linearly with the data size, and running queries multiple times to conduct binary search takes longer. APS thus provides near-optimal performance without tuning overhead, matching or exceeding all baselines.
@@ 8 Discussion
Here we offer a discussion of Quake’s system parameters, how the design extends to new hardware and use-cases with minimal modification.
@@ 8.1 Setting System Parameters
Quake exposes a small number of system parameters, which we fix across all workloads unless otherwise stated. These defaults yield stable performance with minimal tuning and fall into two categories: search parameters and maintenance parameters.
@@
The initial search fraction determines the number of partitions to consider in APS. It has the largest impact on performance. If set too low, APS may not meet the recall target; if too high, the initial scan dominates latency. We default to 5% and reduce this for workloads with many partitions (e.g., >10,000). The number of worker threads is best set to the number of physical cores. Quake scales linearly with thread count until memory bandwidth is saturated (Figure 6). The recompute threshold controls how often APS updates its recall estimate. We set this to 0.01, which avoids unnecessary recomputation with negligible impact on recall (Table 2).
@@
The split/delete threshold $\tau$ sets the minimum predicted latency improvement required to trigger a split or delete. We set $\tau=250$ns. Lower values increase maintenance cost and may cause over-splitting; higher values can allow imbalance to persist. The split penalty parameter $\alpha$ estimates the drop in access frequency after a balanced split. We fix $\alpha=0.5$, which worked well across all benchmarks. If maintenance tuning is needed, we recommend keeping $\alpha$ fixed and adjusting $\tau$ instead. Refinement is controlled by two parameters: the refinement radius $r_{f}$ (number of nearby partitions considered) and the number of refinement iterations. We use one iteration over the 50 nearest partitions. According to our ablation study, disabling refinement reduces maintenance time by 75% but increases query latency and reduces recall (Table LABEL:tab:maintenance_ablation). The window size for access frequency statistics is set equal to the maintenance interval. For example, if maintenance runs every 100,000 queries, the window also spans 100,000. Smaller windows adapt faster but are more volatile.
Most parameters are fixed across workloads. In practice, only the initial search fraction and the maintenance threshold $\tau$ benefit from tuning; however, their defaults are sufficient for the workloads we evaluated.
@@ 8.2 Deployment Considerations
Quake’s design can be extended to support a range of hardware targets and use-cases through minor changes to the cost model and APS logic.
Heterogeneous hardware.
@@
Vector compression.
Using vector compression techniques, such as Product Quantization [15], affect scan cost. Quake can accommodate compression by profiling scan latency over partitions of compressed vectors and updating $\lambda(s)$ in the cost model.
Filters.
Filtered queries can be supported by scaling hit probabilities and access frequencies in APS and the cost model based on estimated selectivity. This will enable Quake to avoid scanning partitions unlikely to contain matching results while preserving recall targets.
Distributed environments.
In a distributed setting, each machine can run APS and maintenance independently on its local partitions using local scan profiles. The cost model can account for partition locality and inter-node variation. A separate load balancer would be required to assign and replicate partitions across machines in a way that minimizes total query cost. We leave the design of such a component to future work.
Concurrency.
@@ Acknowledgments
We’d like to thank our shepherd Nitin Agrawal and the reviewers for their valuable feedback and effort in making this a stronger paper.

### v2

@@ 7.6 Comparison with Early Termination Methods
Table 5 compares early-termination methods on Sift1M, highlighting the tradeoff between query latency, tuning time, and recall, we do not include ground truth generation time in the tuning time for the baselines. APS analytically estimates recall at query time, eliminating offline tuning entirely, while achieving latency within 30% of the oracle across all recall targets. Fixed selects a static $nprobe$ per target via an expensive offline binary search (up to 424s), and SPANN similarly performs a binary search and tunes a centroid-distance threshold; both closely match recall targets but incur higher latency at 90% and 99% recall. LAET trains a per-query prediction model, incurring moderate tuning overhead (81–232s), and matches recall targets with slightly higher latency compared to APS. Auncel is the most similar method to APS, as it aims to analytically estimate recall using partition intersection volumes, however its volume estimation requires calibration, and it is a conservative method, overshooting recall. We tune Auncel by binary searching a geometric parameter ($a$), overshooting recall significantly (up to 8.1 pp) and increasing latency by up to 169% compared to APS. Finally, the Oracle, which scans the minimal amount of partitions per-query, serves as a practical lower bound on achievable latency, though with prohibitively high tuning cost. The tuning overhead of the baselines demonstrates a significant burden for online scenarios where the queries, data, and index change. The tuning burden worsens at scale, where ground truth generation cost grows linearly with the data size, and running queries multiple times to conduct binary search takes longer. APS thus provides near-optimal performance without tuning overhead, matching or exceeding all baselines.
Table 6: Sift10M: Recall and per-level search latency ($\ell_{0}$, $\ell_{1}$, total) for a single-level baseline (L0: 40,000 partitions; L1: 1 partition) and two-level index (L0: 40,000 partitions; L1: 500 partitions), where recall targets are varied at each level. The single-level baseline is the first row of each $\tau_{r}(0)$ block.
$\tau_{r}(0)$ $\tau_{r}(1)$ Recall $\ell_{0}$ (ms) $\ell_{1}$ (ms) Total (ms)
80% 81.2% 2.07 4.85 6.92
80% 74.8% 1.44 0.72 2.16
90% 78.3% 1.56 1.19 2.75
95% 80.1% 1.67 1.69 3.37
99% 81.0% 1.75 2.57 4.33
100% 81.1% 1.82 3.81 5.63
90% 91.3% 2.85 4.89 7.86
80% 84.1% 2.07 0.77 2.84
90% 88.2% 2.26 1.24 3.50
95% 90.1% 2.38 1.72 4.10
99% 91.0% 2.48 2.60 5.08
100% 91.2% 2.62 3.88 6.50
99% 99.0% 4.82 5.5 10.3
80% 91.4% 4.11 0.77 4.88
90% 96.0% 4.59 1.27 5.86
95% 97.7% 4.80 1.75 6.55
99% 98.7% 5.08 2.65 7.74
100% 98.9% 5.28 3.93 9.21
7.7 Multi-Level Recall Estimation
Here we evaluate the effectiveness of APS in a two-level partitioned index by measuring the impact of varying per-level recall targets on overall recall. Table 6 reports results on Sift10M using 40,000 partitions at level $L_{0}$ and 500 partitions at level $L_{1}$. We use an initial search fraction of $f_{M}=1.5\%$ at $L_{0}$ and 25% at $L_{1}$.
We observe that overly aggressive early termination at $L_{1}$ by setting $\tau_{r}(1)$ too low leads to a clear degradation in total recall. For instance, at $\tau_{r}(0)=90\%$, reducing $\tau_{r}(1)$ from 99% to 80% lowers overall recall from 91.0% to 84.1%. This confirms that accurate recall estimation at the upper level is necessary to maintain accurate end-to-end recall estimates. These results justify our design decision in Section 5.1 to fix $\tau_{r}(1)=99\%$ in multi-level configurations so that only the base recall target $\tau_{r}(0)$ needs to be provided.
In addition, the two-level index substantially reduces centroid scanning overhead. The single-level baseline must evaluate distances to all 40,000 centroids per query. In contrast, the two-level configuration performs an approximate search over the centroids. For example, at a recall target of $\tau_{r}(0)=90\%$ with $\tau_{r}(1)=99\%$ the total query latency drops from 7.86ms to 5.08ms, a 35% reduction, driven mainly by the drop in $L_{1}$ latency from 4.89ms to 2.60ms.
@@ 8 Discussion
Here we offer a discussion of Quake’s system parameters, how the design extends to new hardware and use-cases.
@@ 8.1 Setting System Parameters
Quake exposes a few search and maintenance parameters, which we fix across all workloads unless otherwise stated. These defaults give stable performance with minimal tuning.
@@
The initial candidate fraction $f_{M}$ determines the number of partitions to consider in APS. It has the largest impact on performance. If set too low, APS may not meet the recall target; if too high, the initial scan dominates latency. We set this between 1% and 10%. In future work, we aim to remove this parameter entirely. The number of worker threads is best set to the number of physical cores. For large problem sizes, Quake scales linearly with thread count until memory bandwidth is saturated (Figure 6). The recompute threshold controls how often APS updates its recall estimate. We set this to 1%, which avoids unnecessary recomputation with negligible impact on recall (Table 2).
@@
The split/merge threshold $\tau$ sets the minimum predicted latency improvement required to trigger a split or delete. We set $\tau=250$ns. Lower values increase maintenance cost and may cause over-splitting; higher values can allow imbalance to persist. The split access scaling parameter $\alpha$ estimates the drop in access frequency after a split. We fix $\alpha=0.9$, which worked well across all benchmarks. If maintenance tuning is needed, we recommend keeping $\alpha$ fixed and adjusting $\tau$. Refinement is controlled by two parameters: the refinement radius $r_{f}$ (number of nearby partitions considered) and the number of refinement iterations. We use one iteration over the 50 nearest partitions. From our ablation study, disabling refinement reduces maintenance time by 75% but increases query latency and reduces recall (Table LABEL:tab:maintenance_ablation). The window size for access frequency statistics is set equal to the maintenance interval. For example, if maintenance runs every 100,000 queries, the window also spans 100,000 queries. Smaller windows adapt faster but are more volatile.
Most parameters are fixed across workloads. In practice, only the initial candidate fraction $f_{M}$ and the maintenance threshold $\tau$ benefit from tuning; however, their defaults are sufficient for the workloads we evaluated.
@@ 8.2 Deployment Considerations
Quake’s design can be extended to support a range of hardware and use-cases through minor changes to the cost model and APS logic. We reserve these extensions for future work.
Heterogeneous Hardware
@@
Distributed Environments
In a distributed setting, each machine can run APS and maintenance independently on its local partitions. The cost model can account for partition locality and inter-node variation. A separate load balancer would be required to assign and replicate partitions across machines in a way that minimizes total query cost.
Vector Compression
Vector compression techniques, such as Product Quantization [14], reduce scan cost. Quake can support compression by profiling scan latency over partitions of compressed vectors and updating $\lambda(s)$ in the cost model.
Filters
Filtered queries can be supported by scaling per-partition recall probabilities in APS, based on the estimated number of items that pass the filter in each partition. This will enable Quake to avoid scanning partitions unlikely to contain matching results while preserving recall targets.
Concurrency
@@ Acknowledgments
We would like to thank our shepherd, Nitin Agrawal, and the reviewers for their valuable feedback and efforts in making this a stronger paper. This work was supported by NSF grant CNS-2237306, Apple Scholars in AIML PhD Fellowship, and UW-Madison Hilldale Undergraduate Research Fellowship. This work was also supported by DARPA under the grant AIE DARPA-PA-22-01. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views, policies, or endorsements, either expressed or implied, of DARPA or the U.S. Government.
Appendix A Artifact Appendix
Abstract
This artifact provides the experimental setup for Quake to reproduce key results. See the artifact’s README for full details.
Scope
This artifact enables validation of Quake’s core experimental findings, such as those related to Adaptive Partition Scanning (APS), NUMA-aware searching, and maintenance policies. Refer to the artifact’s README for specific experiments.
Contents
The artifact includes Python scripts (e.g., experiment_runner.py, individual experiment run.py files), system installation scripts (install.sh), experiment configurations (configs/), Conda environment files, and the paper PDF. The README details the full directory structure and contents.
Hosting
The artifact is part of the Quake repository in the osdi2025 branch. located at https://github.com/marius-team/quake/tree/osdi2025/test/experiments/osdi2025. Use the latest commit on this branch.
Requirements
Python 3.9+ and Conda are required. Tested on Linux (Ubuntu 22.04). Some experiments (e.g., NUMA evaluations) require a machine with NUMA for meaningful reproduction. The install.sh script lists system-level dependencies. Refer to the README for comprehensive requirements.
A.1 Installation
Follow the detailed installation instructions in the artifact’s README. Options include a comprehensive setup using install.sh or a Quake-only Conda environment setup.
A.2 Experiment Workflow
All experiments are launched via experiment_runner.py from the repository root. Experiments will typically download and prepare required datasets if not found locally. Experiments are run as follows:
python3 -m test.experiments.osdi2025.experiment_runner --experiment kick_the_tires -- config sift1m
After execution, experiments print status updates and save results (e.g., CSV files, plots) to an output directory, as indicated in the console output. The artifact’s README provides the complete command structure, detailed explanations of all parameters (including --output-dir), further examples, and a full summary of available experiments with their specific configurations.